Two small nonconducting spheres have a total charge of 90.0 μCμC .

Two small nonconducting spheres have a total charge of 90.0 μC.

Part A

When placed 1.14 m apart, the force each exerts on the other is 12.0 N and is repulsive. What is the charge on each?

Enter your answers in ascending order separated by comma. Answer in Coulombs.

Q1,Q2 =

Part B

What if the force were attractive?

Enter your answers in ascending order separated by comma.

Q1,Q2 =
Answer #1

Q1 + Q2 = 90 x 10^-6

Force, F = k q1 q2/r^2

12 = 9 x 10^9 x Q1 Q2/1.14^2

Q1 Q2 = 1.73 x 10^-9

Q1 (90 x 10^-6 - Q1) = 1.73 x 10^-9

Q1^2 - 90 x 10^-6 Q1 + 1.73 x 10^-9 = 0

Q1 = 6.21 x 10^-5 C

Q2 = 2.78 x 10^-5 C

B)

Q1 - Q2 = 90 x 10^-6

Force, F = k q1 q2/r^2

12 = 9 x 10^9 x Q1 Q2/1.14^2

Q1 Q2 = 1.73 x 10^-9

Q1 (Q1 - 90 x 10^-6) = 1.73 x 10^-9

Q1^2 - 90 x 10^-6 Q1 - 1.73 x 10^-9 = 0

Q1 = 10.63 x 10^-5 C

Q2 = - 1.63 x 10^-5 C

Comment in case any doubt please rate my answer...